# On the bound states of the Schwarzschild black hole

### v1

@@ I Introduction
Another mapping of the QNM spectrum has been put forward earlier by Mashhoon [22] and applied in more detail in subsequent works [23, 24, 25]. A simple, complex coordinate transformation introduced in Eq. (1) can be used to connect the spectral properties of QNMs with those of the bound states of the “inverted” potential barrier by explicitly transforming the parameters describing the potential. Applying these transformations requires an analytic expression for the bound states. Therefore, Mashhoon’s approach has, until recently [26], been limited to cases where the bound states are known analytically, e.g., in terms of the Pöschl-Teller potential or other exactly solvable potentials. This limited the method to approximate potentials and QNMs with small overtone numbers. In Ref. [27], the inverted potential approach was discussed in connection with the higher-order WKB method, which provides a semi-analytic result, while connections to Seiberg-Witten theory can be found in Ref. [28]. Formal limitations of the bound states to the QNM correspondence are discussed in Ref. [29] and a related application using hyperboloidal methods can be found in Ref. [30]. Note that generalizations of the classical Bohr-Sommerfeld rule have also been used to compute QNMs in Ref. [31].
@@ III Results
Figure 4: Here we compare the bound states of the Regge-Wheeler potential with and without a bump. In the top panel, we show the actual bound states; in the middle panel, we show their absolute difference; and in the bottom panel, the relative difference. We indicate the minimum of the bump $\delta V_{0}$ in the top and middle panel (black solid line) and the ratio of $\delta V_{0}/V_{0}$ in the bottom panel (black solid line). The data is also reported in Table 1 in the supplementary material A.
@@ IV Conclusions
We have also demonstrated that the bound states to QNMs correspondence can provide complementary insights to open problems in the QNM literature, in particular for spectral stability [35, 36, 37, 39, 40, 41]. Because the turning points of the bound-state problem are real, it is also straightforward to relate properties of the potential well with certain bound states. Based on Mashhoon’s identification with QNMs, it is tempting to do the same for the corresponding QNM. In this context, it is worth mentioning that the sensitivity of overtones to modifications of the effective potential close to the horizon has been pointed out recently in Ref. [59, 60]. This finding is qualitatively understandable in the bound state picture presented here.
Since this study is a brief report focusing on the main results, we plan to provide a more complete analysis in a separate work. However, the extension to $n\geq 10$ and $\ell\geq 5$ is straightforward and should provide qualitatively similar findings. Quantitative differences will arise due to the increasing depth of the potential well as a function of $\ell$. Due to isospectrality, the computation for polar perturbations must yield the same conclusions, while those for a test scalar field should be qualitatively similar.

### v2

# Bound States of the Schwarzschild Black Hole

@@ I Introduction
Another mapping of the QNM spectrum was put forward earlier by Mashhoon [22] and applied in more detail in subsequent works [23, 24, 25]. A simple, complex coordinate transformation introduced in Eq. (1) can be used to connect the spectral properties of QNMs with those of the bound states of the “inverted” potential barrier by explicitly transforming the parameters describing the potential. Applying these transformations requires an analytic expression for the bound states. Therefore, Mashhoon’s approach has, until recently [26], been limited to cases where the bound states are known analytically, e.g., in terms of the Pöschl-Teller potential or other exactly solvable potentials. This limited the method to approximate potentials and QNMs with small overtone numbers. In Ref. [27], the inverted potential approach was discussed in connection with the higher-order WKB method, which provides a semi-analytic result, while connections to Seiberg-Witten theory can be found in Ref. [28]. Formal limitations of the bound states to the QNM correspondence are discussed in Ref. [29] and a related application using hyperboloidal methods can be found in Ref. [30]. Note that generalizations of the classical Bohr-Sommerfeld rule have also been used to compute QNMs in Ref. [31].
@@ III Results
Figure 4: Here we compare the bound states of the Regge-Wheeler potential with and without a bump. In the top panel, we show the actual bound states; in the middle panel, we show their absolute difference; and in the bottom panel, the relative difference. We indicate the minimum of the bump $\delta V_{0}$ in the top and middle panel (black solid line) and the ratio of $\delta V_{0}/V_{0}$ in the bottom panel (black solid line). The data are also reported in Table 1 in the supplementary material A.
@@ IV Conclusions
We have also demonstrated that the bound states to QNMs correspondence can provide complementary insights to open problems in the QNM literature, in particular for spectral stability [35, 36, 37, 39, 40, 41]. Because the turning points of the bound-state problem are real, it is also straightforward to relate properties of the potential well with certain bound states. Based on Mashhoon’s identification with QNMs, it is tempting to do the same for the corresponding QNM. In this context, it is worth mentioning that the sensitivity of overtones to modifications of the effective potential close to the horizon has been pointed out recently in Refs. [59, 60]. This finding is qualitatively understandable in the bound state picture presented here.
Since this study is a brief report focusing on the main results, we plan to provide a more complete analysis in a separate work. However, the extension to $n\geq 10$ and $\ell\geq 5$ is straightforward and should provide qualitatively similar findings. Quantitative differences will arise due to the increasing depth of the potential well as a function of $\ell$. Because of isospectrality, the computation for polar perturbations should yield the same conclusions, while those for a test scalar field should be qualitatively similar.
@@ Acknowledgments
Data availability
The data supporting this study’s findings are available within the Letter [65].